An electron emits a photon with a wavelength of 410.0 nanometers (10-9 m).
What is the frequency of this photon?
1
Expert's answer
2020-10-30T06:43:47-0400

f=c/l, where c is the speed of light (greek lambda should be here instead of l)

f=300/4.1•10-7=7.31•108(Hz)